4_._ _Average Variable Costs and the Shutdown Decision

Gob (pronounced Job) runs a banana stand on the oceanfront boardwalk in Newport Beach, CA. He sells frozen bananas at a price of $2.50 per unit. He cannot raise the price without losing all his sales to a nearby competitor. Current sales are 90 frozen bananas per day. He buys bananas at a wholesale price of $0.50 per unit. The cost of freezing bananas is negligible. Operating the banana stand requires a permit from the Business License Office. The permit fee is $1,200 per month, and the fee for this coming month has already been paid. (For the purpose of solving this problem, assume the month has 30 days).

If he were not running the banana stand, Gob could be working as a magician at a local retirement home, earning $50 a day. Alternatively, he could work as an Uber driver, earning $160 a day. Assume Gob is determined to work every single day this month because he wants to buy a new Segway.

a) Compute his average variable cost (in dollars of economic cost per frozen banana). Compare it to the price at which he sells frozen bananas. Should he shut down the banana stand?

A member of the church of the banana offers Gob $50 a day to rent out the banana stand and use it to proselytize and hand out flyers (in which case it cannot be used to sell frozen bananas).

b) Compute his average variable cost (in dollars of economic cost per frozen banana). Compare it to the price at which he sells frozen bananas. What should he do now?

c) How would Gob’s decision be affected by the news that Newport Beach has just passed a tax holiday and will refund all permit fees for the month? Explain your answer.
Answer #1

a) Variable cost = 0.50 *90 = $ 45 per day

Total Fixed cost = 1200/30 = $ 40 per day

Total Cost =$ 85 per day

He is selling banana at 2.50

Total Revenue = $ 225

His profit per banana is 225-85 = 140

His opportunity cost is $ 160 as Uber driver. Therefore, is is suggested that he should shut down.

b) His AVC = 0.50

Contribution = $ 2 per banana

Total Contribution - 180

His opportunity cost is $ 160 as Uber driver. Therefore, is is suggested that he should not shut down. He should continue production.

c) If Newport Beach has just passed a tax holiday then it will refund permit fees and total fixed cost will become zero and therefore he will earn $ 180 per day and hence it is not suggested to him that he should shut down.